OpenStudy (anonymous):

For the given functions, f and g, find the following. f(x) =2x+3 ; g(x)=3x-1 a. f+g b. f-g c. (f)(g) d. f/g

terenzreignz (terenzreignz):

Doesn't look too bad... for instance, the first one, f+g That just means you add the two functions, like so: f+g = f(x) + g(x) = (2x + 3) + (3x - 1) = 2x+3+3x-1= 5x - 2 And that should give you an idea about how the others are to be worked out. Cheers ^_^

OpenStudy (anonymous):

Ok. Thanks!

terenzreignz (terenzreignz):

Post your answers, let's have a look at them (if you like) ^_^